OpenStudy (anonymous):

If a tennis ball has a diameter of approximately 6.7 centimeters, what is its surface area? 282.05 square centimeters 141.03 square centimeters 944.87 square centimeters 564.10 square centimeters

OpenStudy (anonymous):

You need to find the surface area of the sphere, that is given by 4pi*radius squared. Surface area = 4*pi*(6.7/2)^2 = 4*3.142*3.35*3.35 = 141.044 square centimeters

OpenStudy (anonymous):

mubzz you are a lot of help :D